Simplify step by step
13x^2+8x-2x^2+9x

Answer by stanbon(75887)

Simplify step by step
13x^2+8x-2x^2+9x = 11x^2 + 17x = x(11x+17)
---------------------
Cheers,
Stan H.
----------